OpenStudy (anonymous):

Evaluate the function at each specified value of the independent variable and simplify. f(x) = 9 − 2x, x < 0 9, 0 ≤ x < 1 8x + 1, x ≥ 1

OpenStudy (turingtest):

What specified value? There are only intervals... Do they want the endpoints evaluated maybe?

OpenStudy (anonymous):

What values? -1? 0? 1? 2? f(-1) = 9+2 = 11 f(0) = 9 f(0.5) = 9 f(1) = 8 + 1= 9 f(2) = 16 + 1 = 17